# RIS-Enabled NLoS Near-Field Joint Position and Velocity Estimation under User Mobility

### v2

@@ I Introduction
[…]
@@ II-A Geometric Model
The 3D positions are all expressed in the same global reference coordinates system, as follows: $\bm{p}_{b}\in\mathbb{R}^{3\times 1}$ is a vector containing the known BS coordinates, $\bm{p}_{m}\in\mathbb{R}^{3\times 1}$ is a vector holding the known coordinates of the RIS $m$-th element where $m\in\{1,\dots,M\}$ and $r$ is the reference element which is the RIS center, in our case. Also, $\bm{p}\in\mathbb{R}^{3\times 1}$ holds the UE’s coordinates and is expressed as $[p_{x},p_{y},p_{z}]^{\top}$ in the Cartesian coordinates system or as $[\rho,\theta,\phi]^{\top}$ in the spherical coordinates system. Fig. 1 illustrates the scenario, its geometry, and the chosen coordinates axes. Moreover, we denote by $\bm{v}=[v_{x},v_{y},v_{z}]^{\top}\in\mathbb{R}^{3\times 1}$ the UE velocity vector, and by $v_{m}=\bm{v}^{\top}\bm{u}_{m}(\bm{p})$ the radial velocities along the $m$-th RIS element - UE direction. We also define $\bm{u}_{m}(\bm{p})=({\bm{p}_{m}-\bm{p}})/{d_{m}}$, and $d_{m}=\left\lVert\bm{p}_{m}-\bm{p}\right\rVert$.
@@ II-C Problem Description and Maximum-Likelihood Estimator
$(\widehat{\alpha}~{},\widehat{\bm{p}}~{},\widehat{\bm{v}})$ $=\arg\min_{\alpha,\bm{p},\bm{v}}\lVert\bm{y}-\alpha\bm{h}(\bm{p},\bm{v})\rVert^{2},$ (9)
@@ III-A Top-Down Description
The overall method is described in Algorithm 1. The methods commence by assuming a static UE with unknown position, and then iterates between velocity and position estimations until the objective function (9) converges. A final, Quasi-Newton, refinement routine is performed to better estimate the positional, velocity, and gain parameters. The subroutines ${\rm{InitPosGain}}(\bm{y})$ and ${\rm{RefVel}}(\bm{y},\widehat{\bm{v}},\widehat{\bm{p}},\widehat{\alpha})$ will be described in Section III-B, and the subroutine ${\rm{RefPosGain}}(\bm{y},\widehat{\bm{v}},\widehat{\bm{p}},\widehat{\alpha})$ will be detailed in Section III-C.
@@
$(\widehat{\bm{v}},\widehat{\alpha})\leftarrow{\rm{RefVel}}(\bm{y},\widehat{\bm{v}},\widehat{\bm{p}},\widehat{\alpha})~{}.$
@@
$(\widehat{\bm{p}},\widehat{\alpha})\leftarrow{\rm{RefPosGain}}(\bm{y},\widehat{\bm{v}},\widehat{\bm{p}},\widehat{\alpha})~{}.$
@@
9:Perform 6D gradient descent search, starting from $(\widehat{\bm{p}},\widehat{\bm{v}})$.
@@ III-B1 Coarse Estimate
[…]
@@ III-C Velocity Estimate with Known Position
Based on this approximation we then compute the terms in (8) reflecting (33) as $[h_{\ell}(\bm{p},\bm{v}_{0}+\bm{v}_{\delta})]_{m}=\sum_{m=1}^{M}[\bm{w}_{\ell}]_{m}[\bm{a}(\bm{p}_{\ell})]_{m}~{}\approx\nu_{\ell}+j\bm{v}_{\delta}^{\top}\bm{\mu}_{\ell}$ where $\nu_{\ell}=\sum_{m=1}^{M}[\bm{w}_{\ell}]_{m}\exp\big{(}j\beta_{m}\big{)}\exp\big{(}j\bm{v}_{0}^{\top}\bm{\gamma}_{m}\ell T_{s}\big{)},$ and $\bm{\mu}_{\ell}=\sum_{m=1}^{M}[\bm{w}_{\ell}]_{m}\exp\big{(}j\beta_{m}\big{)}\exp\big{(}j\bm{v}_{0}^{\top}\bm{\gamma}_{m}\ell T_{s}\big{)}\bm{\gamma}_{m}\ell T_{s}.$ Using the definition in (7), the vector form of $h_{\ell}(\bm{p},\bm{v}_{0}+\bm{v}_{\delta})$ can be expressed as
@@ IV-D1 Sensitivity to the Speed
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{-3}$$10^{-2}$$10^{-1}$ velocity norm${}\mathrm{[m/s]}$RMSE${}\mathrm{[m]}$PEBGrid SearchCF Position RefinementGlobal Position Refinement
@@ IV-D2 Sensitivity to Uncontrolled Multipath
$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{-3}$$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$$10^{2}$${\mathcal{K}}$RMSE${}\mathrm{[m]}$With MultipathGrid SearchCF Position RefinementWithout MultipathPEBGrid SearchCF Position Refinement
@@ IV-D3 Sensitivity to SNR
We also evaluated the performance of our system w.r.t. $\text{SNR}={|\alpha|^{2}}/({N_{0}n_{f}W})$. Thus, in Fig. 7 we have plotted the PEB performance alongside our devised estimation routines as we change the SNR, all while fixing the RIS-UE distance and UE velocity to $5~{}\mathrm{m}$ and $1~{}\mathrm{m/s}$ respectively. We first notice that the PEB follows an expected trend of linearly decreasing with higher SNR values. On the other hand, Algorithm 2 does not achieve acceptable levels of accuracy at very low SNR values, which means that no localization service could be guaranteed at such conditions. However, as the SNR hits $-20~{}\mathrm{dB}$, the performance of the aforementioned algorithm is boosted to the point where it touches the PEB, and then continuous at that same level even if the SNR is significantly increased. This saturation is due to the finite grid resolution used in this algorithm. Finally, it is noticeable that the CF position refinement algorithm fail completely at very low SNR, but then, also at $-20~{}\mathrm{dB}$ its performance comes back and it follows the trend of the PEB; and of course, since its output is fed as input to the global refinement algorithm, the latter follows the same trend as well.
@@ IV-E Convergence
Finally, we empirically illustrate the average convergence rates (over all the Monte Carlo simulation trials) of both Algorithm 1 (lines 4 through 7) and Algorithm 2 (lines 6 through 8) in Fig. 8(a) and 8(b), respectively, where the RIS-UE distance is set to $2~{}\mathrm{m}$ and the UE’s velocity is set to $1~{}\mathrm{m/s}$. Convergence is considered to be accomplished if the evolution of the respective objective functions (17) and (9) is smaller than an arbitrarily chosen threshold $e\approx 10^{-15}$. Fig. 8(a) shows that (17) converges in just $3$ iterations whereas Fig. 8(a) shows that (9) takes around $20$ iterations to finally converge, which is still acceptable from an algorithmic point-of-view. This practically shows how quick the two algorithmic steps can be in a typical application case. Moreover, Algorithms (3) and 4 converge once the objective functions in (28) and (35) converge to the prespecified threshold, meaning that the error between the received observation and the one reconstructed via the updated position and velocity residuals (respectively) tends to 00.
@@
$10^{0}$$10^{1}$$10^{2}$$10^{-9}$$10^{-8}$$10^{-7}$$10^{-6}$umber of IterationsObjective FunctionObjective Function (9)
@@ Appendix A Derivation of $f_{\ell,m}(\bm{p},\bm{v})$ in (5)
$=\left\lVert\bm{p}_{m}-\bm{p}\right\rVert-\frac{(\bm{p}_{m}-\bm{p})^{\top}}{\left\lVert\bm{p}_{m}-\bm{p}\right\rVert}\bm{v}\ell T_{s}~{},$
$=d_{m}-\bm{u}_{m}^{\top}(\bm{p})\bm{v}\ell T_{s}~{}.$ (47)
@@ Appendix B FF Approximation
[…]
@@ Appendix E PEB & FIM Derivations
FIM$=\frac{2}{P_{n}}\operatorname{Re}\left\{\left(\frac{\partial\mu}{\partial\zeta}\right)^{H}\frac{\partial\mu}{\partial\zeta}\right\}\in\mathbb{R}^{8\times 8},$ (66)
@@
$\frac{\partial\mu}{\partial\zeta}=\begin{bmatrix}\frac{\alpha\omega_{\ell}^{\top}\partial(\bm{p}_{\ell})}{\partial\bm{p}}\vspace{2mm}\\ \frac{\alpha\omega_{\ell}^{\top}\partial(\bm{p}_{\ell})}{\partial\bm{v}}\vspace{2mm}\\ \alpha_{i}\omega_{\ell}^{\top}(\bm{p}_{\ell})\vspace{2mm}\\ \alpha_{r}\omega_{\ell}^{\top}(\bm{p}_{\ell})\end{bmatrix}^{\top}.$ (67)
@@
[…]

### v3

@@ I Introduction
Scalar variables, vectors, and matrices are respectively denoted by lower-case, lower-case bold, and upper-case bold letters (e.g., $a,\bm{a},\bm{A}$) and their indices are denoted by subscripts. Moreover, $\bm{I}_{N}$ and ${{\bm{0}}}_{N}$ represent the identity and the all-zero matrices of size $N\times N$, respectively. The symbols $(.)^{\top}$, $(.)^{*}$ and $(.)^{\mathsf{H}}$ respectively denote transpose, conjugate and hermitian conjugate, the $\mathrm{tr}(\bm{A})$ operator is the trace of matrix $\bm{A}$, and $\text{diag}(\bm{a})$ denotes a diagonal matrix with diagonal elements defined by vector $\bm{a}$. Furthermore, the operators $\|\cdot\|$, $\operatorname{Re}\{\cdot\}$ and $\operatorname{Im}\{\cdot\}$ represent the $l_{2}$-norm real and imaginary operators. Finally, $\mathbb{E}\{\cdot\}$ is the expectation of a random variable with respect to (w.r.t.) its distribution and $\bm{\Pi}^{\perp}_{\bm{A}}={\bm{\mathrm{I}}}-\bm{\Pi}_{\bm{A}}$ the the projection into the null-space of $\bm{A}$ with $\bm{\Pi}_{\bm{A}}=\bm{A}(\bm{A}^{\mathsf{H}}\bm{A})^{-1}\bm{A}^{\mathsf{H}}$.
@@ II-A Geometric Model
The 3D positions are all expressed in the same global reference coordinates system, as follows: $\bm{p}_{b}\in\mathbb{R}^{3\times 1}$ is a vector containing the known BS coordinates, $\bm{p}_{m}\in\mathbb{R}^{3\times 1}$ is a vector holding the known coordinates of the RIS $m$-th element where $m\in\{1,\dots,M\}$ and $r$ is the reference element which is the RIS center, in our case. Also, $\bm{p}\in\mathbb{R}^{3\times 1}$ holds the UE’s coordinates and is expressed as $p_{x},p_{y},p_{z}]^{\top}$ in the Cartesian coordinates system or as $\rho,\theta,\phi]^{\top}$ in the spherical coordinates system. Fig. 1 illustrates the scenario, its geometry, and the chosen coordinates axes. Moreover, we denote by $\bm{v}=[v_{x},v_{y},v_{z}]^{\top}\in\mathbb{R}^{3\times 1}$ the UE velocity vector, and by $v_{m}=\bm{v}^{\top}\bm{u}_{m}(\bm{p})$ the radial velocities along the $m$-th RIS element - UE direction. We also define $\bm{u}_{m}(\bm{p})=({\bm{p}-\bm{p}_{m}})/{d_{m}}$, and $d_{m}=\left\lVert\bm{p}_{m}-\bm{p}\right\rVert$.
@@ II-C Problem Description and Maximum-Likelihood Estimator
[…]
@@ III-A Top-Down Description
[…]
@@
[…]
@@
[…]
@@
[…]
@@ III-B1 Coarse Estimate
where $\bm{k}(\phi,\theta)=-\frac{2\pi}{\lambda}[\sin{\phi}\cos{\theta}~,\hskip 2.84526pt\sin{\phi}\sin{\theta}~,\hskip 2.84526pt\cos{\phi}]^{\top}.$ Leveraging the approximated model in (15), we performed a 2D search over $\theta\in[0,\,2\pi]$ in azimuth and $\phi\in[0,\,\pi/2]$ in elevation to find $\widehat{\theta}$ and $\widehat{\phi}$, respectively:
@@ III-C Velocity Estimate with Known Position
Based on this approximation we then compute the terms in (8) reflecting (33) as $h_{\ell}(\bm{p},\bm{v}_{0}+\bm{v}_{\delta})]_{m}=\sum_{m=1}^{M}[\bm{w}_{\ell}]_{m}[\bm{a}(\bm{p}_{\ell})]_{m}~\approx\nu_{\ell}+j\bm{v}_{\delta}^{\top}\bm{\mu}_{\ell}$ where $\nu_{\ell}=\sum_{m=1}^{M}[\bm{w}_{\ell}]_{m}\exp\big{(}j\beta_{m}\big{)}\exp\big{(}j\bm{v}_{0}^{\top}\bm{\gamma}_{m}\ell T_{s}\big{)},$ and $\bm{\mu}_{\ell}=\sum_{m=1}^{M}[\bm{w}_{\ell}]_{m}\exp\big{(}j\beta_{m}\big{)}\exp\big{(}j\bm{v}_{0}^{\top}\bm{\gamma}_{m}\ell T_{s}\big{)}\bm{\gamma}_{m}\ell T_{s}.$ Using the definition in (7), the vector form of $h_{\ell}(\bm{p},\bm{v}_{0}+\bm{v}_{\delta})$ can be expressed as
@@ IV-D1 Sensitivity to the Speed
$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{-3}$$10^{-2}$$10^{-1}$UE velocity norm[$\mathrm{[m/s]}$RMSE[$\mathrm{[m]}$PEBGrid SearchCF Position RefinementGlobal Position Refinement
@@ IV-D2 Sensitivity to Uncontrolled Multipath
$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{3}$$10^{4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$\mathcal{K}$}RMSE[$\mathrm{[m]}$With MultipathGrid SearchCF Position RefinementWithout MultipathPEBGrid SearchCF Position Refinement
@@ IV-D3 Sensitivity to SNR
We also evaluated the performance of our system w.r.t. $\text{SNR}={|\alpha|^{2}}/({N_{0}n_{f}W})$. Thus, in Fig. 7 we have plotted the PEB performance alongside our devised estimation routines as we change the SNR, all while fixing the RIS-UE distance and UE velocity to $5~\mathrm{m}$ and $1~\mathrm{m/s}$ respectively. We first notice that the PEB follows an expected trend of linearly decreasing with higher SNR values. On the other hand, Algorithm 2 does not achieve acceptable levels of accuracy at very low SNR values, which means that no localization service could be guaranteed at such conditions. However, as the SNR hits $20~\mathrm{dB}$, the performance of the aforementioned algorithm is boosted to the point where it touches the PEB, and then continuous at that same level even if the SNR is significantly increased. This saturation is due to the finite grid resolution used in this algorithm. Finally, it is noticeable that the CF position refinement algorithm fail completely at very low SNR, but then, also at $20~\mathrm{dB}$ its performance comes back and it follows the trend of the PEB; and of course, since its output is fed as input to the global refinement algorithm, the latter follows the same trend as well.
@@ IV-E Convergence
Finally, we empirically illustrate the average convergence rates (over all the Monte Carlo simulation trials) of both Algorithm 1 (lines 4 through 7) and Algorithm 2 (lines 6 through 8) in Fig. 8(a) and 8(b), respectively, where the RIS-UE distance is set to $2~\mathrm{m}$ and the UE’s velocity is set to $1~\mathrm{m/s}$. Convergence is considered to be accomplished if the evolution of the respective objective functions (17) and (9) is smaller than an arbitrarily chosen threshold $e\approx 10^{-15}$. Fig. 8(a) shows that (17) converges in just $3$ iterations whereas Fig. 8(a) shows that (9) takes around $20$ iterations to finally converge, which is still acceptable from an algorithmic point-of-view. This practically shows how quick the two algorithmic steps can be in a typical application case. Moreover, Algorithms (3) and 4 converge once the objective functions in (28) and (35) converge to the prespecified threshold, meaning that the error between the received observation and the one reconstructed via the updated position and velocity residuals (respectively) tends to 0.
@@
$10^{0}$$10^{1}$$10^{2}$$10^{-9}$$10^{-8}$$10^{-7}$$10^{-6}$Number of IterationsObjective FunctionObjective Function (9)
@@ V Conclusion
Acknowledgment
The authors would like to express their sincere gratitude to Dr. Thomas Zemen from the Austrian Institute of Technology (AIT), Austria, for his valuable feedback and insightful suggestions, which helped improve the clarity and accuracy of several expressions in this work.
@@ Appendix A Derivation of $f_{\ell,m}(\bm{p},\bm{v})$ in (5)
$=\left\lVert\bm{p}_{m}-\bm{p}\right\rVert+\frac{(\bm{p}-\bm{p}_{m})^{\top}}{\left\lVert\bm{p}_{m}-\bm{p}\right\rVert}\bm{v}\ell T_{s}~,$
$=d_{m}+\bm{u}_{m}^{\top}(\bm{p})\bm{v}\ell T_{s}~.$ (47)
@@ Appendix B FF Approximation
where $\bm{k}(\phi,\theta)=-\frac{2\pi}{\lambda}[\sin{\phi}\cos{\theta},\hskip 2.84526pt\sin{\phi}\sin{\theta},\hskip 2.84526pt\cos{\phi}]^{\top}$.
@@ Appendix E PEB & FIM Derivations
$\text{\acs{FIM}}=\frac{2}{P_{n}}\operatorname{Re}\left\{\left(\frac{\partial\mu}{\partial\zeta}\right)^{H}\frac{\partial\mu}{\partial\zeta}\right\}\in\mathbb{R}^{8\times 8},$ (66)
@@
$\frac{\partial\mu}{\partial\zeta}=\begin{bmatrix}\frac{\alpha\omega_{\ell}^{\top}\partial(\bm{p}_{\ell})}{\partial\bm{p}}\vskip 5.69054pt\\ \frac{\alpha\omega_{\ell}^{\top}\partial(\bm{p}_{\ell})}{\partial\bm{v}}\vskip 5.69054pt\\ \alpha_{i}\omega_{\ell}^{\top}(\bm{p}_{\ell})\vskip 5.69054pt\\ \alpha_{r}\omega_{\ell}^{\top}(\bm{p}_{\ell})\end{bmatrix}^{\top}$. (67)
@@
PEB $=\sqrt{\text{tr}\left([\acs{FIM}^{-1}]_{1:3,1:3}\right)}$ (68)
VEB $=\sqrt{\text{tr}\left([\acs{FIM}^{-1}]_{4:6,4:6}\right)}.$ (69)